One day a store sold 27 sweatshirts. White ones cost $9.95 and yellow one costs $13.50. In all, $339.65 worth of sweatshirts were sold. How many each color were sold?
Answer by mananth(16946)

x= number of yellow shirts
y= number of white shirts

1.00 x + 1.00 y = 27.00 .............1

13.50 x + 9.95 y = 339.65 .............2
Eliminate y
multiply (1)by -9.95
Multiply (2) by 1.00
-9.95 x -9.95 y = -268.65
13.50 x + 9.95 y = 339.65
Add the two equations
3.55 x = 71.00
/ 3.55
x = 20.00
plug value of x in (1)
1.00 x + 1.00 y = 27.00
20.00 + y = 27.00
y = 27.00 -20.00
y = 7.00
y = 7.00
x= 20 number of yellow shirts
y= 7 number of white shirts
m.ananth@hotmail.ca